# Network-Realised Model Predictive Control Part II: Distributed Constraint Management

### v4

@@ 7.4 Comparison with Alternative Solutions
In order to conclude this numerical example, we present a comparison with several prominent alternative techniques from literature, which tackle the platooning problem in an MPC-based framework. We refer here to the procedures proposed in [plat_MPC_1, plat_MPC_2, plat_MPC_3, plat_MPC_4, plat_MPC_5] and we point out that a major limitation for all of them is the fact that, unlike our proposed solution, no source of disturbance (either from exogenous sources or from model uncertainty) is considered in any of these approaches. Moreover, the control laws presented in [plat_MPC_1, plat_MPC_2, plat_MPC_3] all share in common the fact that they do not explicitly support constraints for the state variables of the platoon’s cars, except for a restrictive use of terminal constraints, which are meant to ensure a number of theoretical guarantees.
Another common feature of the cited works is the fact that the lead cars in [plat_MPC_2, plat_MPC_3, plat_MPC_4] must broadcast information to all other vehicles in the formation, which places an impractical amount of strain on the communication infrastructure of the platoon. In addition to this fact, the property of asymptotic closed-loop stability is retrieved in [plat_MPC_1, plat_MPC_3, plat_MPC_4] by carefully tuning the stage costs of the MPC-based subcontrollers, rather than it being an inbuilt feature, as ensured by the NRF-based first layer (see [part1] and [plutonizare]). This being said, we now address the specific aspects of the five aforementioned techniques:
@@
    One of the key features of the procedure from [plat_MPC_1] is the fact that the above-mentioned terminal constraint requires the computation of a regional average value, evaluated with respect to a number of neighbouring cars. In addition to the demands this places on the communication infrastructure of the platoon, the feasibility of the resulting control problem is noticeably dependent on choosing a large enough prediction horizon. While the overall computational complexity of the optimisation problem to be solved is similar to that of a QP-based formulation in (8), increasing the prediction horizon may have a significant computational impact which can be bypassed in our framework (recall Remark 6.6).
@@
    The main limitation of the technique in [plat_MPC_2] is the fact that the optimisation problem associated with the MPC-based subcontrollers has to be run twice during a car’s command computation routine, with said problem having a comparable cost to a QP-based formulation of our problem in (8). Furthermore, we point out that both the asymptotic and string stability of the closed-loop system are only guaranteed when the platoon’s initial condition satisfies a series of restrictive assumptions.
@@
    In addition to the list of common restrictions detailed above, the solution proposed in [plat_MPC_3] is affected by another limitation; namely, that it can only be initialised when the platoon is already in equilibrium. In spite of these disadvantages, we point out that the procedure’s computational cost is appealingly modest, being roughly analogous to that of our own procedure’s QP-based formulation, since it does not require the double iteration from [plat_MPC_2].
@@
    The approach from [plat_MPC_4] possesses some structural similarity with the one presented in this paper, in the sense that it includes a rudimentary, PID-based lower control layer, while also enabling the use of limited, box-based constraints for the platoon’s states. However, we point out that the lower-level PIDs are not only unable to ensure closed-loop string stability (in contrast to our first layer), but the NRF-based subcontrollers offer many more degrees of freedom, in terms of both sparsity patterns and closed-loop performance (see [part1]). Since [plat_MPC_4] is heavily inspired by the approach presented in [plat_MPC_3], the same computational benefits apply to it as well.
@@
    Finally, the procedure laid out in [plat_MPC_5] proposes a centralised take on the problem and, unsurprisingly, it showcases a high degree of closed-loop performance in the presence of state constraints. In addition to being intractable in the distributed setting, its critical aspect is given by the fact that closed-loop string stability is incompatible with the aforementioned constraints. In contrast to our solution, when safety-critical constraints become active, the technique given in [plat_MPC_5] sacrifices the platoon’s string stability for guaranteed constraint satisfaction.
@@ 8 Conclusions
As a contribution to the synthesis of distributed constraint management techniques, we have shown that it is possible to elegantly decouple structural design constraints (which represent the focal point of the distributed control paradigm) from the time-based constraints which are ubiquitous in MPC-oriented literature. By leveraging a promising class of distributed control strategies [NRF, part1, aug_sparse] and by adapting a series of set-theoretical notions [STMC] to the distributed setting, we obtain a computationally efficient control architecture which provides powerful theoretical guarantees, and whose components can be designed fully offline.
@@
$\mathrm{row}_{j}(w_{r\ell}[k+1])=$ $-a_{j\ell}\,\mathrm{row}_{1}(w_{r\ell}[k])+\mathrm{row}_{j+1}(w_{r\ell}[k])+K_{j\ell}\tiny\begin{bmatrix}u_{f}[k]+\beta_{f}[k]\\ x[k]+\zeta[k]+u_{s1}[k]+\beta_{s1}[k]\end{bmatrix},\,\forall\,j\in\{1:n_{r\ell}-1\},$ (B.1a)
$\hskip-2.84526pt\mathrm{row}_{n_{r\ell}}(w_{r\ell}[k+1])=$ $-a_{n_{r\ell}\ell}\,\mathrm{row}_{1}(w_{r\ell}[k])+K_{n_{r\ell}\ell}\tiny\begin{bmatrix}u_{f}[k]+\beta_{f}[k]\\ x[k]+\zeta[k]+u_{s1}[k]+\beta_{s1}[k]\end{bmatrix}$, (B.1b)
$\mathrm{row}_{\ell}(u_{f}[k])=$ $\ \mathrm{row}_{1}(w_{r\ell}[k]),\,\forall\,k\geq k_{0},\,\ell\in\mathcal{A}_{ui}.$ (B.1c)
$x_{i}[k_{0}+1]=$ $\ C_{xi}\xi_{i}[k_{0}+1]+(\psi_{xi}[k_{0}+1]+\theta_{xi}[k_{0}+1]+\delta_{xi}[k_{0}+1]),\,\psi_{xi}[k_{0}+1]\in\Psi_{xi1},\,\theta_{xi}[k_{0}+1]\in\Theta_{xi1},\,\delta_{xi}[k_{0}+1]\in\Delta_{xi1},$ (B.2a)
$\mathrm{row}_{j}(u_{fi}[k_{0}+1])=$ $\ \mathrm{row}_{j}(C_{ui})\xi_{i}[k_{0}+1]+(\mathrm{row}_{j}(\psi_{ui}[k_{0}+1])+\mathrm{row}_{j}(\theta_{ui}[k_{0}+1])+\mathrm{row}_{j}(\delta_{ui}[k_{0}+1])),\,\mathrm{row}_{j}(\psi_{ui}[k_{0}+1])\in\left(e_{j}^{\top}\Psi_{ui1}\right),$
$\ \mathrm{row}_{j}(\theta_{ui}[k_{0}+1])\in\left(e_{j}^{\top}\Theta_{ui1}\right),\,\mathrm{row}_{j}(\delta_{ui}[k_{0}+1])\in\left(e_{j}^{\top}\Delta_{ui1}\right),\,\forall\,j\in\{1:n_{ui}\}.$ (B.2b)
@@ Appendix A Summary of NRF-based Control
We include here several key definitions and properties associated with the class of control laws from [NRF].
@@
$\mathbf{K}(z)=(I-\mathbf{\Phi}(z))^{-1}\mathbf{\Gamma}(z),\vskip-2.84526pt$
@@
$u[k]=\mathbf{\Phi}(z)\star u[k]+\mathbf{\Gamma}(z)\star x[k].\vskip-8.53581pt$ (58)
@@
In the NRF framework, the problem of imposing sparsity on $\mathbf{K}(z)$ for a distributed implementation (a highly restrictive objective, see [NRF] and [aug_sparse]) is relaxed into ensuring sparsity for its NRF pair $(\mathbf{\Phi}(z),\mathbf{\Gamma}(z))$. The latter goal is far more tractable, due to the manner in which the pair may be computed. To this end, we now introduce the notion of Doubly Coprime Factorisation (DCF) over a set $\mathbb{C}_{g}$, which can be employed as a design parameter.
@@
With respect to the realisation from (1a)-(1c), we define $\mathbf{G}_{u}(z):=(zI_{n_{x}}-A)^{-1}B_{u}$ and we say that any collection of eight TFMs $(\mathbf{N}(z),$ $\mathbf{M}(z),\mathbf{X}(z),\mathbf{Y}(z),\widetilde{\mathbf{N}}(z),$ $\widetilde{\mathbf{M}}(z),\widetilde{\mathbf{X}}(z),\widetilde{\mathbf{Y}}(z))$ is a DCF over $\mathbb{C}_{g}$ of $\mathbf{G}_{u}(z)$ if:
@@
One of the main features of this representation is the fact that, by computing the NRF pair as above, its two terms inherit the sparsity patterns of $\widetilde{\mathbf{Y}}_{\mathbf{Q}}(z)$ and $\widetilde{\mathbf{X}}_{\mathbf{Q}}(z)$. Indeed, since the two TFMs are affine terms of $\mathbf{Q}(z)$, ensuring desired sparsity patterns on these expressions is a particularly tractable problem (see [aug_sparse], for example). A further benefit of employing this control formalism is the fact that all closed-loop maps from exogenous disturbance to the interconnection’s internal signals (recall Remark A.1) are similarly simple expressions of the $\mathbf{Q}(z)$ parameter. For example, the TFM $\mathbf{F}_{\mathbf{Q}}(z)\scriptsize\begin{bmatrix}I_{(n_{x}+n_{u})}&O\end{bmatrix}^{\top}$ from (14) can be expressed as follows
@@
with the other TFMs from (14) being given explicitly in the main result of our companion paper [part1]. Consequently, systemic norm-based procedures (such as the ones proposed in [aug_sparse]) can be used to manipulate these amenable expressions, with the aim of decoupling the areas of the network from (1a)-(1c), as per Remark 4.4.
Appendix B Proofs and Auxiliary Results
@@
The result follows by straightforward algebraic manipulations of the state-space dynamics given in (10). For the sake of clarity, we write these equations out in an explicit manner and we obtain the identities from (B.1a)-(B.1c), which are located at the top of this page.
@@
Moving on, we now prove the result which represents the cornerstone of our constraint management technique.
Proof of Proposition 6.1
The proof is constructive in nature and, to begin with, notice that Assumption A2) enables the local computation of the second-layer command pairs $u_{s1i}[k_{0}]\in\mathcal{U}_{s1i}$ and $u_{s2i}[k_{0}]\in\mathcal{U}_{s2i}$, for each $i\in\{1:N\}$, as in Step 3 of Algorithm 1. Since the algebraic conditions from (13) are satisfied by each of the first layer’s subcontrollers, then there is no need to pass on the values of $u_{s1j}[k_{0}]$, for any $j\in\{1:N\}\setminus\,\mathcal{N}_{i}$, to the $i^{\text{th}}$ subcontroller from the first layer. Therefore, it follows that by communicating and by applying these partial command signals which originate from the second layer, as indicated in Steps 4a and 4b of Algorithm 1, then the implemented command scheme is equivalent to the one depicted in Figure 2. Consequently, the implementation described in Algorithm 1 ensures that the closed-loop dynamics of our control scheme’s first layer are those in (16a)-(16d).
$\xi_{i}[k_{0}+1]=B_{s1i}u_{s1i}[k_{0}]+B_{s2i}u_{s2i}[k_{0}],\vskip-2.84526pt$ (61)
@@
The first conclusion we are able to draw, at this point, is the fact that $x[k_{0}+1]\in\mathcal{X}$, by virtue of (26). Now, in order to retrieve the remaining inclusions given in our result’s statement, notice that all of the assumptions associated with Proposition 5.1 hold. Thus, by applying said result, we get that $\mathrm{row}_{j}(w_{r\ell}[k_{0}+1])\in\mathcal{W}_{\ell j},$ for all $i\in\{1:N\},\,\ell\in\mathcal{A}_{ui}$ and $j\in\{1:n_{r\ell}\}$, along with the fact that $u_{i}[k_{0}]\in\mathcal{U}_{i},\,\forall\,i\in\{1:N\}$. Finally, the latter inclusions yield $u[k_{0}]\in\mathcal{U}$, by way of (26).
We now proceed to employ the arguments made in the proof of Proposition 6.1, in order to obtain guarantees of recursive feasibility for our architecture’s second layer.
@@
The central idea, which underpins the entire proof, is to show that all of the problems described in (8) are guaranteed to be feasible at an arbitrary time instant $k_{0}$, and considering some prediction horizons $T_{i}>0$ and $\overline{T}_{i}\geq 0$. Once this is achieved, the proof boils down to a series of induction-based arguments in which Proposition 6.1 is repeatedly applied, with the aim of ensuring the sought-after inclusions from the result’s statement.
We begin by choosing an arbitrary area index $i\in\{1:N\}$, and by investigating its associated area of the network. Due to Assumption A2), we have the fact that (46) holds for some $\rho_{i}\geq 1$. Thus, for any $\widetilde{\theta}_{it}\in\widetilde{\Theta}_{it}$ and any $t\in\{1:\rho_{i}\}$, there must exist $p_{it}\in\widetilde{\mathcal{P}}_{it}$, $u_{1it}\in\mathcal{U}_{s1i}$ along with $u_{2it}\in\mathcal{U}_{s2i}$ such that the following identities hold
$\scriptsize\begin{array}[]{l}\widetilde{\theta}_{it}=\left(\begin{bmatrix}-C_{xi}\\ -C_{ui}\end{bmatrix}\displaystyle\sum_{j=1}^{t}A_{si}^{j-1}(B_{s1i}u_{1ij}+B_{s2i}u_{1ij})\right)+{p}_{it}.\end{array}\vskip-8.53581pt\normalsize\hskip-5.69054pt$ (63)
@@
$\left\{\scriptsize\begin{array}[]{l}\widetilde{\theta}_{xit}+C_{xi}\displaystyle\sum_{j=1}^{t}A_{si}^{(j-1)}\scriptsize\begin{bmatrix}B_{s1i}&B_{s2i}\end{bmatrix}\hskip-2.84526pt\scriptsize\begin{bmatrix}u_{1ij}\\ u_{2ij}\end{bmatrix}\hskip-2.84526pt={p}_{xit},\\ \widetilde{\theta}_{u\ell t}+\mathrm{row}_{(\ell-\alpha_{ui})}(C_{ui})\displaystyle\sum_{j=1}^{t}A_{si}^{(j-1)}\scriptsize\begin{bmatrix}B_{s1i}&B_{s2i}\end{bmatrix}\hskip-2.84526pt\scriptsize\begin{bmatrix}u_{1ij}\\ u_{2ij}\end{bmatrix}\hskip-2.84526pt={p}_{u\ell t},\end{array}\right$. (64)
for all $t\in\{1:\rho_{i}\}$ and all $\ell\in\mathcal{A}_{ui}$, where $\widetilde{\theta}_{xit}\in\widetilde{\Theta}_{xit}$, $\widetilde{\theta}_{u\ell t}\in\left(e_{(\ell-\alpha_{ui})}^{\top}\widetilde{\Theta}_{uit}\right)$, ${p}_{xit}\in\mathcal{P}_{1it}$ and ${p}_{u\ell t}\in\mathcal{P}_{2\ell t}$. We now employ the expressions deduced in (64) in order to prove the recursive feasibility of the problems formulated in (8). To do so, choose any desired initial time instant $k_{0}$ for the optimisation problem associated with the $i^{\text{th}}$ area index. For all $t\in\{1:\rho_{i}\}$, consider any $\widetilde{\theta}_{xi}[k_{0}+t]\in\widetilde{\Theta}_{xit}$ and any $\widetilde{\theta}_{ui}[k_{0}+t]\in\widetilde{\Theta}_{uit}$ which may arise in the context of the aforementioned optimisation problem, in order to define $\widetilde{\theta}_{xit}:=\widetilde{\theta}_{xi}[k_{0}+t]$ along with $\widetilde{\theta}_{u\ell t}:=\mathrm{row}_{(\ell-\alpha_{ui})}\left(\widetilde{\theta}_{ui}[k_{0}+t]\right)$, with $\ell\in\mathcal{A}_{ui}$. Due to the arguments made in the previous paragraph of this proof, it follows that there must exist some appropriately constrained vectors denoted $u_{1it}$ and $u_{2it}$ along with $p_{xit}$ and $p_{u\ell t}$, with $t\in\{1:\rho_{i}\}$ and $\ell\in\mathcal{A}_{ui}$, such that the identities from (64) hold. Therefore, by setting $T_{i}=\rho_{i}$ and by selecting an arbitrary $\overline{T}_{i}\in\mathbb{N}$, it is always possible to choose the command signal vectors which appear in (8) as $u_{s1i}[k_{0}+t-1]:=u_{1ij}\in\mathcal{U}_{s1i}$ and as $u_{s2i}[k_{0}+t-1]:=u_{2ij}\in\mathcal{U}_{s2i}$, for all $t\in\{1:T_{i}\}$. A direct consequence of all these choices is the fact that
$\xi_{i}[k_{0}+t]=\sum_{j=1}^{t}A_{si}^{(j-1)}(B_{s1i}u_{1ij}+B_{s2i}u_{2ij}),\vskip-8.53581pt$ (65)

### v5

@@ 7.4 Comparison with Alternative Solutions
In order to conclude this numerical example, we present a comparison with several prominent alternative techniques from literature that tackle the platooning problem in an MPC-based framework. We refer here to the procedures proposed in [31, 28, 6, 10, 27], and we point out that a major limitation for all of them is the fact that, unlike our proposed solution, no source of disturbance (either from exogenous sources or from model uncertainty) is considered in any of these approaches. Moreover, the control laws presented in [31, 28, 6] all share in common the fact that they do not explicitly support constraints for the state variables of the platoon’s cars, except for a restrictive use of terminal constraints, which are meant to ensure several theoretical guarantees.
Another common feature of the cited works is that the lead cars in [28, 6, 10] must broadcast information to all other vehicles in the formation, thereby placing an impractical burden on the platoon’s communication infrastructure. In addition to this fact, the property of asymptotic closed-loop stability is retrieved in [31, 6, 10] by carefully tuning the stage costs of the MPC-based subcontrollers, rather than it being an inbuilt feature, as ensured by the NRF-based first layer (see [24] and [20]). In light of these considerations, we now address the specific aspects of the five aforementioned techniques:
@@
    One of the key features of the procedure from [31] is the fact that the above-mentioned terminal constraint requires the computation of a regional average value, evaluated with respect to several neighbouring cars. In addition to the demands this places on the platoon’s communication infrastructure, the feasibility of the resulting control problem is notably dependent on choosing a sufficiently long prediction horizon. While the overall computational complexity of the optimisation problem to be solved is similar to that of a QP-based formulation in (8), increasing the prediction horizon may have a significant computational impact which can be bypassed in our framework (recall Remark 6.6).
@@
    The main limitation of the technique in [28] is the fact that the optimisation problem associated with the MPC-based subcontrollers has to be run twice during a car’s command computation routine, with said problem having a comparable cost to a QP-based formulation of our problem in (8). Furthermore, we note that both the asymptotic and string stability of the closed-loop system are guaranteed only when the platoon’s initial condition satisfies several restrictive assumptions.
@@
    In addition to the list of common restrictions detailed above, the solution proposed in [6] is affected by another limitation; namely, that it can only be initialised when the platoon is already in equilibrium. Despite all of these disadvantages, we point out that the procedure’s computational cost is appealingly modest, roughly comparable to that of our own problem’s QP-based formulation, since it does not require the double iteration from [28].
@@
    The approach from [10] shares some structural similarities with the one presented in this paper, in that it includes a rudimentary PID-based lower control layer, and it also enables the use of limited, box-based constraints on the platoon’s states. However, we point out that the lower-level PIDs are not only unable to ensure closed-loop string stability (in contrast to our first layer), but the NRF-based subcontrollers offer many more degrees of freedom, in terms of both sparsity patterns and closed-loop performance (see [24]). Moreover, since [10] is heavily inspired by the approach presented in [6], the same computational benefits apply to it as well.
@@
    Finally, the procedure outlined in [27] proposes a centralised approach to the problem and, unsurprisingly, demonstrates a high degree of closed-loop performance in the presence of state constraints. In addition to being intractable in the distributed setting, its crucial shortcoming is given by the fact that closed-loop string stability is incompatible with the aforementioned constraints. In contrast to our solution, when safety-critical constraints become active, the technique in [27] sacrifices the platoon’s string stability to guarantee constraint satisfaction.
@@ 8 Conclusions
As a contribution to the synthesis of distributed constraint management techniques, we have shown that it is possible to elegantly decouple structural design constraints (which represent the focal point of the distributed control paradigm) from the time-based constraints which are ubiquitous in MPC-oriented literature. By leveraging a promising class of distributed control strategies [21, 24, 25] and by adapting a series of set-theoretical notions [4] to the distributed setting, we obtain a computationally efficient control architecture which provides powerful theoretical guarantees, and whose components can be designed fully offline.
@@ Appendix A Summary of NRF-based Control
We include here several key definitions and properties associated with the class of control laws from [21].
@@
$\mathbf{K}(z)=(I-\mathbf{\Phi}(z))^{-1}\mathbf{\Gamma}(z),\vskip-5.69054pt$
@@
$u[k]=\mathbf{\Phi}(z)\star u[k]+\mathbf{\Gamma}(z)\star x[k].\vskip-5.69054pt$ (58)
@@
In the NRF framework, the problem of imposing sparsity on $\mathbf{K}(z)$ for a distributed implementation (a highly challenging objective, see [21] and [25]) is relaxed into ensuring sparsity for its NRF pair $(\mathbf{\Phi}(z),\mathbf{\Gamma}(z))$. The latter goal is far more tractable, due to the manner in which the pair may be computed. To this end, we now introduce the notion of Doubly Coprime Factorisation (DCF) over a set $\mathbb{C}_{g}$, which can be employed as a design parameter.
@@
With respect to the realisation from (1a)-(1c), we define $\mathbf{G}_{u}(z):=(zI_{n_{x}}-A)^{-1}B_{u}$ and we say that any collection of eight TFMs $(\mathbf{N}(z),$ $\mathbf{M}(z),\mathbf{X}(z),\mathbf{Y}(z),\widetilde{\mathbf{N}}(z),$ $\widetilde{\mathbf{M}}(z),\widetilde{\mathbf{X}}(z),\widetilde{\mathbf{Y}}(z))$ forms a DCF over $\mathbb{C}_{g}$ of $\mathbf{G}_{u}(z)$ if:
@@
One of the main features of this representation is the fact that, by computing the NRF pair as above, its two terms inherit the sparsity patterns of $\widetilde{\mathbf{Y}}_{\mathbf{Q}}(z)$ and $\widetilde{\mathbf{X}}_{\mathbf{Q}}(z)$. Indeed, since the two TFMs are affine terms of $\mathbf{Q}(z)$, ensuring desired sparsity patterns on these expressions is a particularly tractable problem (see [25], for example). A further benefit of employing this control formalism is the fact that all closed-loop maps from exogenous disturbance to the interconnection’s internal signals (recall Remark A.1) are similarly simple expressions of the $\mathbf{Q}(z)$ parameter. For example, the TFM $\mathbf{F}_{\mathbf{Q}}(z)\scriptsize\begin{bmatrix}I_{(n_{x}+n_{u})}&O\end{bmatrix}^{\top}$ from (14) can be expressed as follows
@@
with the other TFMs from (14) being given explicitly in the main result of our companion paper [24]. Consequently, systemic norm-based procedures (such as the ones proposed in [25]) can be used to manipulate these amenable expressions and to dynamically decouple the areas of the network from (1a)-(1c), as per Remark 4.4.
Appendix B Proofs of the Main Results
@@
The result follows by straightforward algebraic manipulations of the state-space dynamics given in (10). For the sake of clarity, we write these equations out in an explicit manner, and we obtain the identities from (B.1a)-(B.1c), which are located at the top of this page.
@@
Moving on, we now prove the result which serves as the cornerstone of our constraint management technique.
@@
Proof of Proposition 6.1
The proof is constructive in nature and, to begin with, notice that Assumption A2) enables the local computation of the second-layer command pairs $u_{s1i}[k_{0}]\in\mathcal{U}_{s1i}$ and $u_{s2i}[k_{0}]\in\mathcal{U}_{s2i}$, for each $i\in\{1:N\}$, as in Step 3 of Algorithm 1. Since the algebraic conditions from (13) are satisfied by each of the first layer’s subcontrollers, then there is no need to pass on the values of $u_{s1j}[k_{0}]$, for any $j\in\{1:N\}\setminus\,\mathcal{N}_{i}$, to the $i^{\text{th}}$ subcontroller from the first layer. Therefore, it follows that by communicating and by applying these partial command signals which originate from the second layer, as indicated in Steps 4a and 4b of Algorithm 1, the implemented command scheme is equivalent to the one shown in Figure 2. Consequently, the implementation described in Algorithm 1 ensures the fact that the closed-loop dynamics of our control scheme’s first layer are those in (16a)-(16d).
$\xi_{i}[k_{0}+1]=B_{s1i}u_{s1i}[k_{0}]+B_{s2i}u_{s2i}[k_{0}],\vskip-8.53581pt$ (62)
@@
The first conclusion we can draw, at this point, is the fact that $x[k_{0}+1]\in\mathcal{X}$, by virtue of (26). Now, to retrieve the remaining inclusions given in our result’s statement, notice that all of the assumptions associated with Proposition 5.1 hold. Thus, by applying said result, we get that $\mathrm{row}_{j}(w_{r\ell}[k_{0}+1])\in\mathcal{W}_{\ell j},$ for all $i\in\{1:N\},\,\ell\in\mathcal{A}_{ui}$ and $j\in\{1:n_{r\ell}\}$, along with the fact that $u_{i}[k_{0}]\in\mathcal{U}_{i},\,\forall\,i\in\{1:N\}$. Finally, the latter inclusions yield the fact that $u[k_{0}]\in\mathcal{U}$, by way of (26).
We now proceed to employ the arguments made in the proof of Proposition 6.1 to obtain guarantees of recursive feasibility for our architecture’s second layer.
@@
The central idea, which underpins the entire proof of this result, is to show that all of the problems described in (8) are guaranteed to be feasible at an arbitrary time instant $k_{0}$, and considering some prediction horizons $T_{i}>0$ and $\overline{T}_{i}\geq 0$. Once these guarantees are obtained, the proof boils down to a series of induction-based arguments in which Proposition 6.1 is repeatedly applied to ensure the sought-after inclusions from the result’s statement.
We begin by choosing an arbitrary index $i\in\{1:N\}$, and by investigating its associated area of the network. Due to Assumption A2), we have that (46) holds for some $\rho_{i}\geq 1$. Thus, for any $\widetilde{\theta}_{it}\in\widetilde{\Theta}_{it}$ and any $t\in\{1:\rho_{i}\}$, there must exist $p_{it}\in\widetilde{\mathcal{P}}_{it}$, $u_{1it}\in\mathcal{U}_{s1i}$ along with $u_{2it}\in\mathcal{U}_{s2i}$ such that the following identities hold:
$\begin{array}[]{l}\widetilde{\theta}_{it}=\left(\begin{bmatrix}-C_{xi}\\ -C_{ui}\end{bmatrix}\displaystyle\sum_{j=1}^{t}A_{si}^{j-1}(B_{s1i}u_{1ij}+B_{s2i}u_{1ij})\right)+{p}_{it}.\end{array}$ (64)
@@
[…]
for all $t\in\{1:\rho_{i}\}$ and all $\ell\in\mathcal{A}_{ui}$, where $\widetilde{\theta}_{xit}\in\widetilde{\Theta}_{xit}$, $\widetilde{\theta}_{u\ell t}\in\left(e_{(\ell-\alpha_{ui})}^{\top}\widetilde{\Theta}_{uit}\right)$, ${p}_{xit}\in\mathcal{P}_{1it}$ and ${p}_{u\ell t}\in\mathcal{P}_{2\ell t}$. We now employ the expressions deduced in (65) to prove the recursive feasibility of the problems stated in (8).
To do so, choose any desired initial time instant $k_{0}$ for the optimisation problem associated with the $i^{\text{th}}$ area index. For all $t\in\{1:\rho_{i}\}$, consider any $\widetilde{\theta}_{xi}[k_{0}+t]\in\widetilde{\Theta}_{xit}$ and any $\widetilde{\theta}_{ui}[k_{0}+t]\in\widetilde{\Theta}_{uit}$ which may arise in the context of the aforementioned optimisation problem, to define $\widetilde{\theta}_{xit}:=\widetilde{\theta}_{xi}[k_{0}+t]$ and $\widetilde{\theta}_{u\ell t}:=\mathrm{row}_{(\ell-\alpha_{ui})}\left(\widetilde{\theta}_{ui}[k_{0}+t]\right)$, where $\ell\in\mathcal{A}_{ui}$. Due to the arguments made in the previous paragraph of this proof, it follows that there must exist some appropriately constrained vectors denoted $u_{1it}$ and $u_{2it}$ along with $p_{xit}$ and $p_{u\ell t}$, with $t\in\{1:\rho_{i}\}$ and $\ell\in\mathcal{A}_{ui}$, such that the identities from (65) hold. Thus, by setting $T_{i}=\rho_{i}$ and by selecting an arbitrary $\overline{T}_{i}\in\mathbb{N}$, it is always possible to choose the command signal vectors appearing in (8) as $u_{s1i}[k_{0}+t-1]:=u_{1ij}\in\mathcal{U}_{s1i}$ and as $u_{s2i}[k_{0}+t-1]:=u_{2ij}\in\mathcal{U}_{s2i}$, for all $t\in\{1:T_{i}\}$.
A direct consequence of all these choices is the fact that
$\xi_{i}[k_{0}+t]=\sum_{j=1}^{t}A_{si}^{(j-1)}(B_{s1i}u_{1ij}+B_{s2i}u_{2ij}),\vskip-11.38109pt$ (66)